# Incremental Learning for Robot Shared Autonomy

### v1

@@ Abstract
Shared autonomy holds promise for improving the usability and accessibility of assistive robotic arms, but current methods often rely on costly expert demonstrations and lack the ability to adapt post-deployment. This paper introduces ILSA, an Incrementally Learned Shared Autonomy framework that continually improves its assistive control policy through repeated user interactions. ILSA leverages synthetic kinematic trajectories for initial pretraining, reducing the need for expert demonstrations, and then incrementally finetunes its policy after each manipulation interaction, with mechanisms to balance new knowledge acquisition with existing knowledge retention during incremental learning. We validate ILSA for complex long-horizon tasks through a comprehensive ablation study and a user study with 20 participants, demonstrating its effectiveness and robustness in both quantitative performance and user-reported qualitative metrics. Code and videos are available at https://ilsa-robo.github.io/.
@@ I INTRODUCTION
Assistive devices such as wheelchair-mounted robotic arms have the potential to benefit millions of individuals with physical disabilities [36] and enhance their independence [5, 2, 27]. However, controlling assistive robotic arms via the current standard of teleoperation is challenging, especially for those with limited mobility. Shared autonomy, where human operators and autonomous systems share decision-making, can make these devices more intuitive and effective for complex tasks, enhancing performance and safety [33, 13].
Recent shared autonomy systems have begun employing imitation learning, which learns an assistive control policy from human demonstrations [23, 20, 16, 8, 39]. However, training these models often requires substantial “expert” demonstrations, which are costly and difficult to scale [14, 29, 40]. Additionally, most shared autonomy and imitation learning implementations focus on creating robust systems before deployment but often remain static once in use, lacking mechanisms for continuous improvement. Without the ability to adapt post-deployment, these systems are limited by their initial performance.
@@
Figure 1: We introduce an Incrementally Learned Shared Autonomy (ILSA) system that is initially pretrained on synthetic kinematic trajectories and then continuously refined as the user repeatedly interacts with it. Left: During the initial trials, ILSA lacks the knowledge required to avoid collisions, leading to task objects colliding with obstacles. Right: With only a few interactions, by the fourth trials, ILSA has learned effective collision avoidance from past user interactions, assisting the user in completing tasks more smoothly.
Incremental learning [37, 12] offers a promising solution to these challenges by enabling neural networks to continuously learn and improve from non-stationary streams of data. This paradigm is particularly well-suited for shared control systems, which are used repeatedly and generate continual interaction data.
Building on these insights, we introduce an Incrementally Learned Shared Autonomy (ILSA) framework (Fig. 2), which improves its generated assistive actions after each user interaction. ILSA begins with a pretraining phase using synthetic kinematic trajectories generated based on simplistic and intuitive rules as described in Section III-B, thereby eliminating the need for actual expert demonstrations. This pretraining phase enables ILSA to provide a basic level of assistance the first time a user uses it. After pretraining, ILSA is incrementally finetuned with each interaction. We design the finetuning process of ILSA to balance learning from new data and retaining existing knowledge. (Refer to Fig. 1 for an illustration of ILSA’s learning progression.) We demonstrate the effectiveness of the designs through a quantitative ablation study.
Figure 2: Our Incrementally Learned Shared Autonomy (ILSA) framework. $a_{m,t}$ is the intermediate output action without user guidance, generated by $\mathcal{F^{\prime}}_{\theta^{\prime}}(s_{t})$. $a_{f,t}$ is the final output action with user guidance, generated by ′′$\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}(s_{t},a_{u,t})$. $\mathcal{F^{\prime}}_{\theta^{\prime}}$ and ′′$\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}$ are subsets of the overall action generation model $\mathcal{F}_{\theta}(s_{t},a_{u,t})$, with blue and green annotations indicating their components, respectively. Their roles are detailed in Section III-C. $L_{demo,m}$, $L_{demo,f}$, $L_{order}$, and $L_{direc}$ are loss terms designed to balance task completion and the user’s sense of control, as defined in Section III-C. The red components highlight the training designs used during the incremental learning phase.
To formally evaluate ILSA, we conducted a user study where 20 participants used a Kinova robotic arm to perform two complex, long-horizon tasks. Study results support our hypothesis that (1) ILSA supports faster task completion and easier manipulator control as compared to pure teleoperation; and (2) ILSA improves task performance over time as a user repeatedly performs a task, in contrast to a static shared autonomy method.
@@
  • We propose ILSA, a novel Incrementally Learned Shared Autonomy framework that improves a shared control policy through continual user interactions without extensive expert demonstrations.
  • Through a quantitative ablation study, we investigate how ILSA balances learning from new data while retaining existing knowledge during finetuning.
  • We conducted a human study, demonstrating that ILSA rapidly enhances task performance over time, significantly outperforming teleoperation and non-incremental shared autonomy.
@@ II RELATED WORK
II-A Shared Autonomy
Despite varying definitions and settings, most shared autonomy approaches aim to map human inputs—often suboptimal due to limited knowledge or experience [31, 39], or constrained to a lower dimension than the robot’s action space [23, 15]—into robot movements that successfully accomplish the task. Regardless of the specific approach, existing research primarily focuses on developing a robust system before deployment, which often requires a complex expert demonstration collection process, especially in recent learning-based methods [14, 29, 20, 8, 39]. This focus likely stems from the perspective that human operators in shared autonomy systems are typically non-experts in need of assistance. However, we contend that even non-expert human operators can contribute positively to the task, and their intelligence should be harnessed. Thus, we propose ILSA, a framework that continuously evolves and improves through repeated user interaction.
@@ II-B Human-in-the-loop Machine Learning
Another related area within full autonomy is human-in-the-loop machine learning [9, 1], where an autonomous agent learns from human interactions to improve performance. Various forms of human feedback, such as demonstrations, preferences, rankings [4], interventions [17, 25, 35], and scalar-valued critiques [24, 38], are integrated into these processes.
Among these works, interactive imitation learning [6, 28, 3], which incorporates human interventions into imitation learning, is particularly relevant to our work. However, interactive imitation learning methods aim to develop fully autonomous systems using expert interventions and still require extensive expert demonstrations to train on before deployment [17, 25, 34, 7, 22, 21]. In contrast, we focus on a shared autonomy system that assists non-expert operators, pretrained on synthetic kinematic trajectories and finetuned through interactions with users. Moreover, existing methods often need numerous policy deployments and substantial human feedback before each update, limiting practical use. Our method updates the policy after each use by the operator, balancing learning from new data while retaining existing knowledge.
II-C Incremental Learning
Incremental learning, sometimes referred to interchangeably with continual learning and lifelong learning, aims to enable data-driven models to incrementally learn from non-stationary streams of data [12]. A core challenge in incremental learning is mitigating a phenomenon known as catastrophic forgetting [26, 10], where new learning can cause previously acquired capabilities to be lost.
The field of incremental learning has seen significant growth in recent years, with applications in various areas such as medical diagnosis [19], autonomous driving [32], and financial forecasting [30]. However, there is minimal work that incorporates incremental learning into robot shared autonomy. Our approach seeks to bridge this gap by leveraging incremental learning to continuously enhance the shared autonomy system’s performance over time through ongoing interactions with human operators.
@@ III-A Problem Statement
Our goal is to identify a function $\mathcal{F}_{\theta}(s_{t},a_{u,t})$: $\mathcal{S}\times\mathcal{A}_{u}\rightarrow\mathcal{A}_{r}$ that maps the task state and user action to a robot action that is both task-optimal and aligned with the user’s intent. While $\mathcal{F}_{\theta}$ may be imperfect initially, the objective is to improve it over time as the user interacts with the shared autonomy system.
III-B Synthetic Kinematic Trajectories Generation
We use synthetic kinematic trajectories to warm-start ILSA, reducing the need for expert demonstrations. To generate these trajectories, we divide multi-step tasks into distinct subtasks. For each subtask, we randomly sample task object locations, calculate the corresponding target end-effector positions, and use linear interpolation between initial and target end-effector positions to create trajectories.
For example, in a cereal-pouring task divided into three subtasks—(1) grasp the cereal container, (2) move the container next to the bowl, and (3) pour the cereal—we randomly sample locations for the cereal container and bowl within the robot’s workspace. Based on these object locations, we compute the robot end-effector’s target positions for each subtask. Sequences of end-effector positions are then generated via linear interpolation between initial and target end-effector positions, forming task states $s_{t}$. Differences between successive positions yield the robot actions $a_{r,t}$. Note that these kinematic trajectories may be imperfect and might not account for factors like obstacle avoidance. However, warm-starting ILSA with diverse trajectories supports broader generalization capabilities without needing to collect expert demonstrations for each new task.
These synthetic kinematic trajectories yield tuples $(s_{t},a_{r,t})$.
@@ III-B Synthetic Kinematic Trajectories Generation
To train ILSA on kinematic trajectories, we also generate synthetic user actions $a_{u,t}$ to mimic real user inputs. We design $a_{u,t}$ as a three-dimensional vector, with each value representing the user’s intended end-effector movement in the $x$, $y$, and $z$ translational dimensions. To generate such $a_{u,t}$, for each dimension $i\in\{x,y,z\}$, if the target end-effector position in dimension $i$ is greater than the current position in the trajectory, we sample $a_{u,t,i}$ to be in $[0,1]$, otherwise we sample $a_{u,t,i}$ in $[-1,0]$. In our real-world experiments, the user provides $a_{u,t}$ using two joysticks on an Xbox controller. One joystick operates two axes of end effector motion (e.g., $x$- and $y$-axes in translation mode) and the other joystick operates the third axis.
In summary, we generate trajectories each consisting of a series of $(s_{t},a_{u,t},a_{r,t})$ triplets. We generate 50 trajectories for each task.

### v3

@@ Abstract
Shared autonomy holds promise for improving the usability and accessibility of assistive robotic arms, but current methods often rely on costly expert demonstrations and remain static after pretraining, limiting their ability to handle real-world variations. Even with extensive training data, unforeseen challenges—especially those that fundamentally alter task dynamics, such as unexpected obstacles or spatial constraints—can cause assistive policies to break down, leading to ineffective or unreliable assistance. To address this, we propose ILSA, an Incrementally Learned Shared Autonomy framework that continuously refines its assistive policy through user interactions, adapting to real-world challenges beyond the scope of pre-collected data. At the core of ILSA is a structured fine-tuning mechanism that enables continual improvement with each interaction by effectively integrating limited new interaction data while preserving prior knowledge, ensuring a balance between adaptation and generalization. A user study with 20 participants demonstrates ILSA’s effectiveness, showing faster task completion and improved user experience compared to static alternatives. Code and videos are available at https://ilsa-robo.github.io/.
@@ I INTRODUCTION
Assistive devices such as wheelchair-mounted robotic arms have the potential to benefit millions of individuals with physical disabilities and enhance their independence [5, 2]. Shared autonomy, where human operators and autonomous systems share decision-making, has emerged as a promising way to make these devices more intuitive and effective [27, 10].
Recent shared autonomy systems have begun employing imitation learning, which learns an assistive control policy from human demonstrations to enhance generalization across task environments [19, 16, 13, 7, 30]. However, collecting large-scale demonstrations is costly and difficult to scale [11, 25, 31]. Moreover, even with extensive training data, it is impractical to cover the full range of real-world scenarios. As a result, assistive policies trained solely on pre-collected data may struggle when encountering previously unseen situations during deployment. This limitation becomes particularly critical when the training data fails to anticipate specific real-world challenges that fundamentally alter the task dynamics, such as the presence of obstacles or spatial constraints. For example, in a cereal pouring task where the system needs to move a cereal container next to a bowl to pour, as shown in the upper two sub-figures of Fig. 1, even a well-designed training set may not account for all possible real-world variations, such as the presence of obstacles between the two objects. If such obstacles exist in deployment, a policy trained without exposure to these cases may struggle to adapt, leading to suboptimal or even infeasible behavior.
@@
Figure 1: We introduce ILSA, an Incrementally Learned Shared Autonomy system that continuously refines its assistive policy through user interaction, enabling adaptation to real-world challenges beyond the scope of pre-collected training data. Left: In early trials, ILSA struggles with collision avoidance due to obstacles or spatial constraints not accounted for in training data. Right: With only a few interactions, by the fourth trials, ILSA has learned effective collision avoidance from past user interactions, assisting the user in completing tasks more smoothly.
To address the inability of static policies to adapt to unexpected real-world challenges that alter task execution, we introduce an Incrementally Learned Shared Autonomy (ILSA) framework (Fig. 2) that continuously fine-tunes its assistive policy as the user interacts with the system. A key technical contribution of ILSA is its structured fine-tuning mechanism, which updates the assistive policy after each user interaction, ensuring enhanced assistance in every subsequent use. This frequent adaptation is particularly challenging due to limited deployment data, which can affect model stability and generalizability. To address this, our fine-tuning mechanism consists of three key components: (1) Corrected Trajectory Supervision, which prevents the model from reinforcing suboptimal actions by generating corrected trajectories from user interactions and using them for fine-tuning; (2) Layered Supervision, which integrates new interaction data with pretraining knowledge to balance adaptation and retention, without requiring manual data reweighting; and (3) Partial Model Update, which selectively fine-tunes critical components to enhance adaptation while maintaining generalization and robustness. These designs collectively enable continual adaptation while preserving previously acquired knowledge, ensuring ILSA retains the generalization ability gained from pretraining. The effectiveness of our fine-tuning approach is validated through a quantitative ablation study.
While our fine-tuning mechanism enables ILSA to continuously adapt to user interactions, the initial pretraining phase remains essential for ensuring effective first-time use. Instead of relying on costly demonstrations that still cannot capture all real-world challenges, we initialize ILSA with a lightweight set of simulated kinematic trajectories, generated through simple rule-based procedures. This allows for efficient task initialization while minimizing training costs.
To formally evaluate ILSA, we conducted a user study with 20 participants using a Kinova robotic arm to perform two complex, long-horizon tasks. These tasks were specifically designed to introduce challenges that were not accounted for in the training data, such as the presence of obstacles or spatial constraints, highlighting scenarios where pre-collected demonstrations may fail to generalize. Crucially, we note that these challenges are not a consequence of using simulated data but rather reflect the broader issue that even expert demonstrations cannot account for every possible variation. Our user study results show that: (1) ILSA enables faster task completion and more intuitive manipulator control compared to pure teleoperation; (2) ILSA effectively adapts to real-world challenges beyond the scope of training data, continuously improving task performance over time, in contrast to a static shared autonomy method.
@@
  • We propose ILSA, an Incrementally Learned Shared Autonomy framework that continuously improves its assistive policy through user interaction, enabling adaptation to real-world challenges beyond the scope of pre-collected data.
  • We design a structured fine-tuning mechanism that enables ILSA to improve its assistive policy after each user interaction, ensuring continual improvement while balancing adaptation and generalization ability. We validate its effectiveness through a quantitative ablation study.
  • We conducted a human study, demonstrating that ILSA rapidly enhances task performance over time, significantly outperforming teleoperation and non-incremental shared autonomy in real-world tasks with previously unseen challenges.
@@ II RELATED WORK
II-A Shared Autonomy with Continual Adaptation
In recent years, a limited number of shared autonomy works explored frameworks that refine assistive policies through repeated user interactions. Two notable works in this direction are Confidence-Aware Shared Autonomy (CASA) [32] and Shared Autonomy across Repeated Interactions (SARI) [12]. Both methods aim to improve assistive policies over time by expanding the robot’s skill set through learning new tasks or goals from user demonstrations. Specifically, CASA estimates confidence in a user’s intent to determine whether a task is known or novel. If recognized, it blends user and robot control; otherwise, it relinquishes control entirely, allowing the user to demonstrate the task. CASA then applies inverse reinforcement learning (IRL) to infer a cost function and incorporate the new task into its skill set. Similarly, SARI automates familiar tasks based on prior interactions while deferring control to the user for novel ones. It periodically retrains its policy on collected trajectories, enabling continual refinement over time.
While these approaches enhance assistance across repeated interactions, they focus on expanding the robot’s skill set by learning new tasks rather than adapting within an existing one. Once a task is learned, they assume it remains unchanged. In contrast, our work addresses a different but equally critical challenge: even within a single task, real-world variations—especially those that fundamentally alter task dynamics, such as unforeseen obstacles or spatial constraints—can significantly impact task execution, requiring continuous adaptation.
Furthermore, CASA and SARI require multiple user demonstrations before updating and retrain periodically on aggregated interaction data, which can lead to long periods of suboptimal performance before policy updates, limiting the user experience. In contrast, we aim to fine-tune the assistive policy after each interaction, enabling continuous adaptation in real-time. Achieving this, however, presents key challenges including ensuring stability and generalization when updating with limited new data. To address this, we focus on effectively integrating new interaction data while preserving pretraining knowledge, striking a balance between adaptation and retention. While prior shared autonomy frameworks have not explicitly explored this direction, we introduce a structured fine-tuning mechanism specifically designed for this challenge. Through our key designs, ILSA ensures that each update contributes to improved assistance, effectively enabling adaptation to real-world challenges beyond the scope of pre-collected data with minimal interaction overhead.
Figure 2: Our Incrementally Learned Shared Autonomy (ILSA) framework. The base action generation model $\mathcal{F}_{\theta}(s_{t},a_{u,t})$ consists of an intermediate action predictor $\mathcal{F^{\prime}}_{\theta^{\prime}}(s{t})$ (blue) and a final action predictor ′′$\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}(s_{t},a_{u,t})$ (green). At the core of ILSA is its structured fine-tuning mechanism (highlighted in red), which incorporates three key components: Corrected Trajectory Supervision, Layered Supervision, and Partial Model Update. These designs enable ILSA to continuously improve with every user interaction while preserving generalization ability, effectively adapting to real-world challenges beyond pre-collected training data.
@@ II-B Human-in-the-loop Machine Learning
Another related area is human-in-the-loop machine learning [8, 1], where autonomous agents refine their policies through human interactions. A variety of human feedback modalities—such as demonstrations, preferences, rankings[4], interventions [14, 21, 28], and scalar-valued critiques [20, 29]—have been explored to enhance autonomous decision-making. Among these approaches, interactive imitation learning [6, 24, 3], which incorporates human interventions into imitation learning, is particularly relevant to our work. A notable example is DAgger (Dataset Aggregation)[26] and its variants, such as EnsembleDAgger[23] and HG-DAgger [14], which address covariate shift in imitation learning by iteratively collecting expert interventions on novel states encountered during deployment, improving policy robustness.
While effective for fully autonomous systems, these methods rely on expert interventions, whereas our work focuses on shared autonomy, where real-time assistance is needed for non-expert users. Furthermore, while human-in-the-loop learning has been shown to improve policy performance, existing approaches primarily evaluate adaptation to common task variations—such as differences in object positions or trajectories—rather than situations that fundamentally alter task execution, like unforeseen obstacles or spatial constraints. On the contrary, ILSA is motivated by the need for adaptation in these challenging scenarios and demonstrates its ability to handle them in our experiments.
Additionally, human-in-the-loop frameworks typically require multiple policy deployments and substantial expert feedback before updating the model, leading to extended periods of suboptimal performance. In contrast, ILSA specifically addresses the challenge of per-interaction adaptation, introducing a structured fine-tuning mechanism that seamlessly integrates new interaction data while preserving pretraining knowledge. This design enables stable and generalizable adaptation even with limited data, ensuring continuous improvement after each user interaction.
@@ III-A Problem Statement
Our goal is to identify a function $\mathcal{F}_{\theta}(s_{t},a_{u,t})$: $\mathcal{S}\times\mathcal{A}_{u}\rightarrow\mathcal{A}_{r}$ that maps the task state and user action to a robot action that is both task-optimal and aligned with the user’s intent. In our work, we model $\mathcal{F}_{\theta}$ using a robot action generation model (Fig. 2). Unlike traditional methods that solely rely on static pre-collected demonstrations to train this model, we adopt a two-stage learning process:
1) Pretraining with lightweight, simulated kinematic trajectories to provide an initial assistive policy.
2) Incremental adaptation through real-world user interactions to continually refine the policy and address unforeseen deployment challenges.
In the following sections, we first describe our simulated trajectory generation process (Section III-B), then introduce our robot action generation model (Section III-C), which serves as the basis of ILSA’s assistive policy. Finally, we present our structured fine-tuning mechanism (Section III-D), the key technical contribution that addresses challenges such as learning from sparse deployment data and maintaining generalization ability, enabling ILSA to continuously adapt to user interactions.
III-B Simulated Kinematic Trajectories Generation
Instead of relying on costly expert demonstrations, which are still unlikely to cover all real-world variations, we use a simple, rule-based approach to generate simulated kinematic trajectories to pretrain ILSA.
To generate these trajectories, we decompose multi-step tasks into distinct subtasks based on predefined task objectives. Each subtask corresponds to a critical transition in task execution, such as object grasping, transport, or tool manipulation. For instance, in a cereal-pouring task consisting of three subtasks—(1) grasping the cereal container, (2) moving it next to the bowl, and (3) pouring—the container and bowl positions are randomly sampled within the workspace, and the corresponding end-effector poses (positions and orientations) necessary to complete each subtask are computed. The robot’s motion is then computed via linear interpolation between initial and target end-effector poses, forming state-action tuples $(s_{t},a_{r,t})$. Here, $s_{t}$ includes the full task state, encompassing the robot’s end-effector pose as well as the poses of all relevant objects in the scene.
@@ III-B Simulated Kinematic Trajectories Generation
To further simulate user interaction, we generate synthetic user actions $a_{u,t}$ to mimic user inputs. The user action is defined as a 3D vector representing intended end-effector movements in the $x$, $y$, and $z$ directions. Each dimension corresponds to a normalized velocity command: if the target position in dimension $i$ is greater than the current position, we sample $a_{u,t,i}$ from $[0,1]$; otherwise, we sample from $[-1,0]$. In real-world deployment, these inputs are provided by a user via a dual-joystick Xbox controller, where one joystick operates two translational axes (e.g., $x$ and $y$), and the other controls the third axis ($z$).
We generate 50 trajectories per task to ensure sufficient diversity for pretraining. Each trajectory consists of a sequence of $(s_{t},a_{u,t},a_{r,t})$ triplets. This lightweight, automated data generation strategy ensures an effective initialization. The simplicity of the trajectory generation process also facilitates easy transferability across different manipulation tasks without requiring complex environment-specific modeling.